OpenStudy (anonymous):

*I don't want the flat out answer i want help * Which of the following exponential functions goes through the points (1, 12) and (2, 36)? f(x) = 4(3)−x f(x) = 3(4)−x f(x) = 3(4)x f(x) = 4(3)x

OpenStudy (amistre64):

trial and error are a mathmatikers best tools

OpenStudy (amistre64):

also, your notation as it posted is not useful. a^x indicates an exponent of x

OpenStudy (anonymous):

Oh yeah , i keep forgetting to do that . Sorry !

OpenStudy (amistre64):

the longhand version, is just to setup 2 equations in 2 unknowns that match the stated points; let y = ab^x 12 = ab^1 36 = ab^2 a = 12/b = 36/b^2 12/36 = b/b^2 .... 3 = b

OpenStudy (anonymous):

Thank you so much !

OpenStudy (amistre64):

i prolly messed that up in there somewhere ... but im outa practice by a few decades :)

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

Thanks :)